OpenStudy (anonymous):

rewrite using commutative property.....8+12

OpenStudy (anonymous):

8 + 12 = 12 + 8

mathslover (mathslover):

remember that : \[\large{a+b=b+a}\] --> Commutative property .. hence 8 + 12 can also be written 12 + 8